OpenStudy (anonymous):

If a ball dropped from a tower reaches the ground after 3.5 seconds, what is the height of the tower?

OpenStudy (anonymous):

Assuming "dropped" means "released" (i.e. initial velocity 0) we can find the height from the following equation: \[h = 0.5*g*t^2\] plugging in the values we get: \[h = 0.5*10*3.5^2 = 61.25 (m)\]

OpenStudy (anonymous):

answer choices are a)56.00 meters b)60.03 meters c)62.08 meters d)62.50 meters

OpenStudy (anonymous):

plugging in g = 9.8 instead of 10: \[h = 0.5*9.8*3.5^2 = 60.025 \approx 60.03\] that's b. I probably had an arithmetic mistake up there.

OpenStudy (anonymous):

nope, simply the difference between g = 10 and 9.8...

OpenStudy (anonymous):

A feather is dropped onto the surface of the moon. How far will the feather have fallen if it reaches the surface in 9.00 seconds? Given: g on moon = -1.63 meters/second2 A)14.0 meters B)14.7 meters C)16.00 meters D)66.0 meters

OpenStudy (anonymous):

Again, assuming it was dropped from rest and by the formula: \[y = y_0 + v_0*t + 0.5*g*t^2\] with: \[y=? ; y_0 = 0 ; v_0 = 0; g = 1.63(m/\sec^2)\] (I took plus instead of minus since I'm taking the positive direction as being down) we get: \[y = 0.5*1.63*9^2 = 66.015 (m)\] That's D.